# Does Graph Prompt Work? A Data Operation Perspective with Theoretical Analysis

### v1

@@ Abstract
In recent years, graph prompting has emerged as a promising research direction, enabling the learning of additional tokens or subgraphs appended to the original graphs without requiring retraining of pre-trained graph models across various applications. This novel paradigm, shifting from the traditional “pre-training and fine-tuning” to “pre-training and prompting” has shown significant empirical success in simulating graph data operations, with applications ranging from recommendation systems to biological networks and graph transferring. However, despite its potential, the theoretical underpinnings of graph prompting remain underexplored, raising critical questions about its fundamental effectiveness. The lack of rigorous theoretical proof of why and how much it works is more like a “dark cloud” over the graph prompt area to go further. To fill this gap, this paper introduces a theoretical framework that rigorously analyzes graph prompting from a data operation perspective. Our contributions are threefold: First, we provide a formal guarantee theorem, demonstrating graph prompts’ capacity to approximate graph transformation operators, effectively linking upstream and downstream tasks. Second, we derive upper bounds on the error of these data operations by graph prompts for a single graph and extend this discussion to batches of graphs, which are common in graph model training. Third, we analyze the distribution of data operation errors, extending our theoretical findings from linear graph models (e.g., GCN) to non-linear graph models (e.g., GAT). Extensive experiments support our theoretical results and confirm the practical implications of these guarantees.
@@ 1 Introduction
Inspired by the success of prompting techniques in natural language processing (NLP), there has been a growing interest in adapting similar ideas to graph data through “pre-training and prompting”. Graph prompts (Sun et al., 2023b) modify the input graphs by adding learnable tokens or subgraphs, enabling the pre-trained GNN to better align with the requirements of downstream tasks without tuning the model parameters. Many empirical works (Sun et al., 2022; Liu et al., 2023; Tan et al., 2023; Huang et al., 2023; Ma et al., 2023; Chen et al., 2024) have found that graph prompting can reduce computational overhead, preserve the generality of the pre-trained model, and allow for seamless application across multiple tasks to achieve better expressive capability than the traditional paradigm.
@@
Recently, some studies (Fang et al., 2024; Sun et al., 2023a) have realized that the reason why graph prompts work may relate to their capability in simulating various data operations like deleting/adding nodes/edges, changing node features, and even removing subgraphs. This makes graph prompts stand out from their counterpart in the NLP area and inspires many empirical applications like recommendation systems (Yang et al., 2023; 2024; Wang et al., 2024), biological networks (Diao et al., 2022; Gao et al., 2024), transferring knowledge across different graph domains (Guo et al., 2023; Zhu et al., 2024b; Zhao et al., 2024), and more (Jin et al., 2024). Unfortunately, despite these promising results, the theoretical basis of graph prompting remains underexplored. Existing works primarily rely on empirical validation and lack rigorous theoretical analysis to explain why graph prompts are effective and how they can be systematically designed. This gap is just like a “dark could” over the graph prompt area, raising critical questions about their broader applications and the development of more advanced methods that could leverage their full potential.
@@
In light of these limitations, this paper provides a comprehensive theoretical framework for graph prompting from a data operation perspective. First, we establish rigorous guarantee theorems that demonstrate the underlying reason why graph prompts work is their capacity to simulate various graph data operations, and our theorems further answer why such capacity can make the pre-trained model meet new task requirements without retraining. Second, we derive upper bounds on the error introduced by graph prompts when simulating these data operations. We analyze this error for individual graphs and extend our discussion to batches of graphs, which is crucial for understanding the scalability and generalization of graph prompts in practical scenarios where models are usually trained on multiple graphs. Third, we explore the distribution of the data operation error and extend our theoretical findings from linear graph models, such as Graph Convolutional Networks (GCNs), to non-linear models like Graph Attention Networks (GATs). This extension demonstrates the robustness of our theoretical framework across different GNN architectures and provides insights into how non-linearity affects the effectiveness of graph prompts. We conduct extensive experiments to confirm our theoretical findings. By offering such a solid theoretical foundation for graph prompting, our work advances the understanding of how and why graph prompts work, guides for designing more effective prompting techniques, and empowers researchers and practitioners to leverage them with greater confidence in various applications.
@@ 2 Background
GPF and All-in-One. Current graph prompt designs, as described in the review by Sun et al. (2023b), can be primarily categorized into two types: prompt as token vectors added to node features, and prompt as additional graph inserted to the original graph. In the rest of this paper, we focus on two representative frameworks: GPF (Fang et al., 2022) as an example of adding extra prompt vectors, and All-in-One (Sun et al., 2023a) as an example of adding prompt subgraphs. Our choice is motivated by our interest in simulating graph operations at a theoretical level. The GPF and All-in-One frameworks provide the most fundamental approaches among these methods. The rest graph prompt designs can usually be treated as their special cases or natural extensions.
@@
All-in-One offers the prompt as a graph format by defining prompt tokens, token structures and inserting patterns. Let $\Omega\in\mathbb{R}^{k\times F}$ be the learnable matrix corresponding to $K$ prompt tokens. Let $\mathbf{A}_{in}\in\{0,1\}^{k\times k}$ indicate token structures where $A_{ij}=1$ means there is an inner link connecting the $i$-th and the $j$-th tokens and vice versa. $A_{in}$ can be calculated by the inner product of these tokens. Similarly, the inserting pattern tells us how to connect each token to the original graph nodes, which is denoted by a cross matrix $\mathbf{A}_{cro}\in\{0,1\}^{k\times N}$. Then the graph prompt changes the original graph $G$ to $G_{\omega}=(\mathcal{V}_{\omega},\mathcal{E}_{\omega},\mathbf{X},\mathbf{\Omega},\mathbf{A}_{\omega})$ where $\mathcal{V}_{\omega}$ is the collection of the original nodes and the token nodes; $\mathcal{E}_{\omega}$ includes the original edges, inner links among tokens and the cross links between tokens and the original nodes; $\mathbf{A}_{\omega}$ is the collection of $\mathbf{A}$,$\mathbf{A}_{in}$, and $\mathbf{A}_{cro}$.
@@ 4.1 Upper Bound of the Error on A Single Graph
Given the All-in-One-like prompt graph $SG_{\omega}$, if a GCN model $F_{\theta}$ does not have any non-linear transformations, or has non-linear layers but the model’s weight matrix is row full-rank, then there exists an optimal $\omega$ for any input graph $G$ such that $P_{\omega}(G)\in B_{G}$.
@@
For the cases where weight matrices are not full-rank, we have found that the error value ($\epsilon$) of the extended bridge set ($\epsilon-B_{G}$), into which the prompting framework can map $G$, is positively correlated with the distance of the graph model parameter matrix from being full-rank. However, a consistent upper bound does exist for the error $\epsilon$ of the extended bridge set when the matrix rank is determined. This means that even in some extreme cases, graph prompt learning can still guarantee a certain level of performance without experiencing extremely unexpectedly poor results:
@@
For a GCN model $F_{\theta}$, assume at least one layer’s parameter matrix is not full rank, for GPF or All-in-One prompt, there exists an upper bound of $\epsilon$ such that for any input graph $G$, there exists an optimal $\omega$ where $P_{\omega}(G)\in\epsilon\text{-}B_{G}$, with $\epsilon\leq\mu(\theta^{*})\cdot\lambda(G)$, where $\mu(\theta^{*})$ and $\lambda(G)$ correspond to the model and graph $G$, respectively.
Theorem 5 is proved in Appendix A.3.3, where the upper bound of the error $\epsilon$ can be further expressed as follows:
$\mu(\theta^{*})\lambda(G)=\sin(\Phi/2)\|C(G)\|$ (4)
The upper bound has two terms, one part is related to the model because $\sin(\Phi/2)$ can be treated as a measurement of the model’s expressiveness, the details of which can be seen in Appendix A.3.3. The other part is related to $\|C(G)\|$. Since the graph prompt aims to approximate $G$ with $P_{\omega}(G_{ori})$, that means the error is also related to the prompt design. In this way, we in theory confirm the intuitional findings by Sun et al. (2023a) as mentioned in our motivation section.
Figure 1: Error w.r.t matrices rank.
The upper bound of $\epsilon$ given in Theorem 5 reveals the potential distortion of $B_{G}$’s shape when the matrix is not full-rank and the model’s expressive power is insufficient. This can lead to an increased distance between $B_{G}$ and the transformation domain $D_{P}(G)$ (see this term in section 3.3) of GPF or All-in-One prompts. To confirm this judgment, we conducted a quantitative analysis using numerical methods for the case of non-full rank matrices. We make the graph model weight matrix be not full-rank (with rank $n-r$ where $n$ is the full rank number and $r=0,\cdots n$) and use the distance between the embedding vectors of $P_{\omega}(G)$ and $C(G)$ as the loss function. We applied stochastic gradient descent with a learning rate of 0.0001 to optimize $\omega$ until convergence. To avoid local optima, we repeated the process multiple times with different initializations for each graph. The results are shown in Figure 1, where the vertical axis is the empirical max $\epsilon$ of the extended $B_{G}$ and the horizontal axis is $r$ where $r=0,1,2,\cdots,9,10$, making the matrix rank become $n-r$. From Figure 1 we can find that with the increase of $r$, the rank of the model matrix becomes lower, making the model expressiveness worse and then leading to a larger error bound. Besides, more advanced graph prompts (e.g., GPF-Plus, All-in-One, and All-in-One-Plus) generally have a lower bound than the naive one (e.g., GPF).
@@ 4.2 Extend the Error Bound Discussion to A Batch of Graphs
Here for the $M$ graphs in $\mathcal{G}$, we first construct an optimal solution matrix according to function $C$, thus have: $S=[C(G_{1}),\ldots,C(G_{M})]$. Then $V=S^{\top}S\in\mathbb{R}^{M\times M}$ denotes the correlation matrix of downstream solutions upon such graph dataset. The eigenvalues of $V$ sorted by the descending order can be denoted as $\{\lambda_{1},\cdots,\lambda_{M}\}$. Then the upper bound $\epsilon^{*}$ can be treated as the mean square over the smallest $M-k$ eigenvalues. In practice, the eigenvalues of $V$ in datasets often exhibit an exponential decay (Johnstone, 2001). This explains the rapid decrease in error rate as $k$ increases, proving that prompts are not only effective but also efficient. With the increasing number of graphs $M$ in the dataset, the largest $k$ ($k\ll M$) eigenvalues can almost explain most of the matrix, which means using small-scaled prompt tokens can achieve reasonably accurate results. This finding is also consistent with many existing empirical researches (Liu et al., 2023; Sun et al., 2023a; Wang et al., 2024; Zhu et al., 2024a; Zi et al., 2024).
@@ 4.3 Value Distribution of the Data Operation Error with Graph Prompt
Consider a graph model with Leaky-Relu as their activation function, we denote the graph embedding residual vector as $\bm{\beta}\in\mathbb{R}^{1\times F}$ where $F$ is the graph embedding dimensions. Each entry of $\bm{\beta}$ is related to the bridge graph embedding, graph prompt, and the original graph, which is mathematically defined as follows:
$[\bm{\beta}]_{j}=\sum_{i}-\alpha^{k}w_{i}v_{ij}$ (6)
Figure 2: real $\epsilon$ distribution and fitted curves
where $\alpha$ is the parameter of Leaky-Relu; $v_{ij}$ is the $j$-th element in the $i$-th node embedding of the bridge graph $G_{bri}$; $w_{i}$ is the weight of the $i$-th node for graph pooling (e.g., summation graph pooling means for any node $i$ in $G_{bri}$, we have $w_{i}=1$); $k$ can be either 00 or $-1$ and depends on the original graph, graph prompt, and the bridge graph, the details of which can be seen in Appendix 5. Here, we assume that $\bm{\beta}$ should follow an i.i.d. normal distribution with mean 0 and variance $c$ (we carefully discuss the rationality of this assumption in Appendix 5): $\bm{\beta}\sim\mathcal{N}(0,cI_{n})$ where $I_{n}$ is the $n\times n$ identity matrix and $c$ is a positive constant. Then we theoretically find that $\epsilon$ conforms to the Chi distribution ($\chi_{k}$):
@@
4.4 Extend the Discussion from Linear Models to Non-linear Models
@@
Let $F_{\theta}$ be a GAT model. If any layer of the model has a full row rank parameter matrix, then for the All-in-One prompting framework, for any input graph G, there exists an optimal $\omega$ such that $P_{\omega}(G)\in B_{G}$. When the parameter matrix is not full rank, there is an upper bound $\mu(\theta)\cdot\lambda(G)$ making $P_{\omega}(G)\in\epsilon\text{-}B_{G}$, $\epsilon\leq\mu(\theta)\cdot\lambda(G)$.
@@ 5.1 Experimental Settings
(a) GCN
(b) GAT
Figure 3: convergence rate analysis
Model Settings: We utilize two GNN frameworks: GCN (representing linear models) and GAT (representing non-linear models). We limit our experiments to these two models as other models follow similar patterns. Unless otherwise specified, we use a 3-layer GNN with Leaky-ReLU activation function and feature dimension $F=25$. For full-rank matrix studies, we ensure each layer’s matrix is full-rank (selected after pre-training). For non-full-rank matrix studies, we set the rank loss to 5 by default. The default ReadOut method is mean pooling.
Training: In the graph prompting training process, we perform gradient descent on the parameters $\omega$ of the graph prompt $P_{\omega}$ using the Adam optimizer. We use a learning rate of $1\times 10^{-4}$ and weight decay of $5\times 10^{-5}$. We implement an early stopping mechanism with a maximum of 2,000 epochs by default. Our loss function is defined as $\|F_{\theta^{*}}(G_{p})-C(G)\|$ for single-graph tasks, and $\sqrt{\sum_{G\in\mathcal{G}}\|F_{\theta^{*}}(G_{p})-C(G)\|^{2}/M}$ for multi-graph tasks where $G_{p}$ is the combined graph with $G$ and graph prompt, and $C(G)$ means an optimal function to the downstream task, which is not accessible without a specific task. Since the ultimate purpose of graph prompting is to approximate graph operation, we here treat $C(\cdot)$ as various graph data permutations such as adding/deleting nodes, adding/deleting/changing edges, and transforming features of a given graph $G$. Then we wish to see how well the graph prompt reaches $C(G)$ by manipulating graph data with a graph prompt. For more detailed experimental settings, please check in the Appendix C.

### v2

@@ Abstract
In recent years, graph prompting has emerged as a promising research direction, enabling the learning of additional tokens or subgraphs appended to the original graphs without requiring retraining of pre-trained graph models across various applications. This novel paradigm, shifting from the traditional “pre-training and fine-tuning” to “pre-training and prompting” has shown significant empirical success in simulating graph data operations, with applications ranging from recommendation systems to biological networks and graph transferring. However, despite its potential, the theoretical underpinnings of graph prompting remain underexplored, raising critical questions about its fundamental effectiveness. The lack of rigorous theoretical proof of why and how much it works is more like a “dark cloud” over the graph prompt area to go further. To fill this gap, this paper introduces a theoretical framework that rigorously analyzes graph prompting from a data operation perspective. Our contributions are threefold: First, we provide a formal guarantee theorem, demonstrating graph prompts’ capacity to approximate graph transformation operators, effectively linking upstream and downstream tasks. Second, we derive upper bounds on the error of these data operations by graph prompts for a single graph and extend this discussion to batches of graphs, which are common in graph model training. Third, we analyze the distribution of data operation errors, extending our theoretical findings from linear graph aggregations (e.g., GCN) to non-linear graph aggregations (e.g., GAT). Extensive experiments support our theoretical results and confirm the practical implications of these guarantees.
Graph Prompt, Graph Neural Network
@@ 1 Introduction
Inspired by the success of prompting techniques in natural language processing (NLP), there has been a growing interest in adapting similar ideas to graph data through “pre-training and prompting”. Graph prompts (Sun et al., 2023b) modify the input graphs by adding learnable tokens or subgraphs, enabling the pre-trained GNN to better align with the requirements of downstream tasks without tuning the model parameters. Many empirical works (Sun et al., 2022; Liu et al., 2023; Tan et al., 2023; Huang et al., 2023; Ma et al., 2023) have found that graph prompting can reduce computational overhead, preserve the generality of the pre-trained model, and allow for seamless application across multiple tasks to achieve better expressive capability than the traditional paradigm.
@@
Recently, some studies (Fang et al., 2024; Sun et al., 2023a) have realized that the reason why graph prompts work may relate to their capability in simulating various data operations like deleting/adding nodes/edges, changing node features, and even removing subgraphs. This makes graph prompts stand out from their counterpart in the NLP area and inspires many empirical applications like recommendation systems (Yang et al., 2023, 2024), biological networks (Diao et al., 2022), transferring knowledge across different graph domains (Guo et al., 2023; Zhu et al., 2024b), and more. Unfortunately, despite these promising results, the theoretical basis of graph prompting remains underexplored. Existing works primarily rely on empirical validation and lack rigorous theoretical analysis to explain why graph prompts are effective and how they can be systematically designed. This gap is just like a “dark could” over the graph prompt area, raising critical questions about their broader applications and the development of more advanced methods that could leverage their full potential.
@@
In light of these limitations, this paper provides a comprehensive theoretical framework for graph prompting from a data operation perspective. First, we establish rigorous guarantee theorems that demonstrate the underlying reason why graph prompts work is their capacity to simulate various graph data operations, and our theorems further answer why such capacity can make the pre-trained model meet new task requirements without retraining. Second, we derive upper bounds on the error introduced by graph prompts when simulating these data operations. We analyze this error for individual graphs and extend our discussion to batches of graphs, which is crucial for understanding the scalability and generalization of graph prompts in practical scenarios where models are usually trained on multiple graphs. Third, we explore the distribution of the data operation error and extend our theoretical findings from linear graph aggregations, such as Graph Convolutional Networks (GCNs), to non-linear aggregations like Graph Attention Networks (GATs). This extension demonstrates the robustness of our theoretical framework across different GNN architectures and provides insights into how non-linearity affects the effectiveness of graph prompts. We conduct extensive experiments to confirm our theoretical findings. By offering such a solid theoretical foundation for graph prompting, our work advances the understanding of how and why graph prompts work, guides for designing more effective prompting techniques, and empowers researchers and practitioners to leverage them with greater confidence in various applications.
@@ 2 Background
GPF and All-in-One. Current graph prompt designs, as described in the review by Sun et al. (2023b), can be primarily categorized into two types: prompt as token vectors added to node features, and prompt as additional graph inserted to the original graph. In the rest of this paper, we focus on two representative frameworks: GPF (Fang et al., 2022) as an example of adding extra prompt vectors, and All-in-One (Sun et al., 2023a) as an example of adding prompt subgraphs. Our choice is motivated by our interest in simulating graph operations at a theoretical level. The GPF and All-in-One frameworks provide the most fundamental approaches among these methods. The rest graph prompt designs can usually be treated as their special cases or natural extensions (Sun et al., 2023b).
@@
All-in-One offers the prompt as a graph format by defining prompt tokens, token structures and inserting patterns. Let $\Omega\in\mathbb{R}^{k\times F}$ be the learnable matrix corresponding to $K$ prompt tokens. Let $\mathbf{A}_{in}\in\{0,1\}^{k\times k}$ indicate token structures where $A_{ij}=1$ means there is an inner link connecting the $i$-th and the $j$-th tokens and vice versa. $\mathbf{A}_{in}$ can be calculated by the inner product of these tokens. Similarly, the inserting pattern tells us how to connect each token to the original graph nodes, which is denoted by a cross matrix $\mathbf{A}_{cro}\in\{0,1\}^{k\times N}$. Then the graph prompt changes the original graph $G$ to $G_{\omega}=(\mathcal{V}_{\omega},\mathcal{E}_{\omega},\mathbf{X},\mathbf{\Omega},\mathbf{A}_{\omega})$ where $\mathcal{V}_{\omega}$ is the collection of the original nodes and the token nodes; $\mathcal{E}_{\omega}$ includes the original edges, inner links among tokens and the cross links between tokens and the original nodes; $\mathbf{A}_{\omega}$ is the collection of $\mathbf{A}$,$\mathbf{A}_{in}$, and $\mathbf{A}_{cro}$.
@@ 4.1 Upper Bound of the Error on A Single Graph
Given the All-in-One-like prompt graph $SG_{\omega}$ (a subgraph containing prompt tokens and token structures), if a GCN model $F_{\theta}$ does not have any non-linear transformations, or has non-linear layers but the model’s weight matrix is row full-rank, then there exists an optimal $\omega$ for any input graph $G$ such that $P_{\omega}(G)\in B_{G}$.
@@
For the cases where weight matrices are not full-rank, we have found that the error value ($\epsilon$) of the extended bridge set ($\epsilon\text{-}B_{G}$), into which the prompting framework can map $G$, is positively correlated with the distance of the graph model parameter matrix from being full-rank. However, a consistent upper bound does exist for the error $\epsilon$ of the extended bridge set when the matrix rank is determined. This means that even in some extreme cases, graph prompt learning can still guarantee a certain level of performance without experiencing extremely unexpectedly poor results:
@@
For a GCN model $F_{\theta}$, assume at least one layer’s parameter matrix is not full rank, for GPF or All-in-One prompt, there exists an upper bound of $\epsilon$ such that for any input graph $G$, there exists an optimal $\omega$ where $P_{\omega}(G)\in\epsilon\text{-}B_{G}$, with $\epsilon\leq\zeta(\theta)\cdot\kappa(G)$, where $\zeta(\theta)$ is an implicit function corresponding to the model and $\kappa(G)$ is an implicit function corresponding to graph $G$, denoting the two parts of the error boundaries.
We recommend the readers go further to see the details of Theorem 5 as proved in Appendix A.3.3, where the upper bound has two terms: one part ($\zeta(\theta)$) is only related to the model itself and increases as the model’s expressive power decreases, such as increasing with decreasing rank; The other part ($\kappa(G)$) is related to the complexity of the solution space of downstream task on the given graph. Note that this is the upper bound of the error, not the error itself. In practice, we use the graph prompt to approximate such solution space, which means a better prompt design makes the real error even less than this upper bound. In this way, we in theory confirm the intuitional findings by Sun et al. (2023a) as mentioned in our motivation section.
@@ 4.2 Extend the Error Bound Discussion to A Batch of Graphs
Here for the $M$ graphs in $\mathcal{G}$, we first construct an optimal solution matrix according to function $C$, thus have: $S=[C(G_{1}),\ldots,C(G_{M})]$. Then $V=S^{\top}S\in\mathbb{R}^{M\times M}$ denotes the correlation matrix of downstream solutions upon such graph dataset. The eigenvalues of $V$ sorted by the descending order can be denoted as $\{\lambda_{1},\cdots,\lambda_{M}\}$. Then the upper bound $\epsilon^{*}$ can be treated as the mean square over the smallest $M-k$ eigenvalues. In practice, the eigenvalues of $V$ in datasets often exhibit an exponential decay (Johnstone, 2001). This explains the rapid decrease in error rate as $k$ increases, proving that prompts are not only effective but also efficient. With the increasing number of graphs $M$ in the dataset, the largest $k$ ($k\ll M$) eigenvalues can almost explain most of the matrix, which means using small-scaled prompt tokens can achieve reasonably accurate results. This finding is also consistent with many existing empirical researches (Liu et al., 2023; Sun et al., 2023a; Wang et al., 2024; Zhu et al., 2024a).
@@ 4.3 Value Distribution of the Data Operation Error with Graph Prompt
Consider a graph model with Leaky-Relu as their activation function, we denote the graph embedding residual vector as $\boldsymbol{\beta}\in\mathbb{R}^{1\times F}$ where $F$ is the graph embedding dimensions. Each entry of $\boldsymbol{\beta}$ is related to the bridge graph embedding, graph prompt, and the original graph, which is mathematically defined as follows:
$[\boldsymbol{\beta}]_{j}=\sum_{i}(-\alpha^{k}w_{i}v_{ij})$ (5)
Figure 1: Real $\epsilon$ distribution and fitted curves.
where $\alpha$ is the parameter of Leaky-Relu; $v_{ij}$ is the $j$-th element in the $i$-th node embedding of the bridge graph $G_{bri}$; $w_{i}$ is the weight of the $i$-th node for graph pooling (e.g., summation graph pooling means for any node $i$ in $G_{bri}$, we have $w_{i}=1$); $k$ can be either 00 or $-1$ and depends on the original graph, graph prompt, and the bridge graph, the details of which can be seen in Appendix 5. Here, we assume that $\boldsymbol{\beta}$ should follow an i.i.d. normal distribution with mean 0 and variance $c$ (we carefully discuss the rationality of this assumption in Appendix 5): $\boldsymbol{\beta}\sim\mathcal{N}(0,cI_{n})$ where $I_{n}$ is the $n\times n$ identity matrix and $c$ is a positive constant. Then we theoretically find that $\epsilon$ conforms to the Chi distribution ($\chi_{k}$):
@@
4.4 Extend the Discussion from Linear to Non-linear
@@
Let $F_{\theta}$ be a GAT model. If any layer of the model has a full row rank parameter matrix, then for the All-in-One prompting framework, for any input graph G, there exists an optimal $\omega$ such that $P_{\omega}(G)\in B_{G}$. When the parameter matrix is not full rank, there is an upper bound $\zeta(\theta)\cdot\kappa(G)$ making $P_{\omega}(G)\in\epsilon\text{-}B_{G}$, $\epsilon\leq\zeta(\theta)\cdot\kappa(G)$.
@@ 5.1 Experimental Settings
Model Settings: We utilize two GNN frameworks: GCN (representing linear models) and GAT (representing non-linear models). We limit our experiments to these two models as other models follow similar patterns. Unless otherwise specified, we use a 3-layer GNN with Leaky-ReLU activation function and feature dimension $F=25$. For full-rank matrix studies, we ensure each layer’s matrix is full-rank (selected after pre-training). For non-full-rank matrix studies, we set the rank loss to 5 by default. The default ReadOut method is mean pooling. For more detailed experimental settings, please check in the Appendix C. We open our testing code at https://github.com/qunzhongwang/dgpw
Figure 2: Convergence rate analysis. GCN (left) and GAT (right).